OpenStudy (anonymous):

how do you solve 4sin5θcos5θ in terms of single trigonometric function?

OpenStudy (callisto):

Use sin2x = 2sinxcosx Put x= 5θ in this case 4sin5θcos5θ = 2 ( 2sin5θcos5θ) = ...?

OpenStudy (anonymous):

why did you put 2(sin5θcos5θ)?

OpenStudy (callisto):

For you to observe easier and apply that identity.

OpenStudy (anonymous):

so it becomes 2(sin5θcos5θ)= 2sin10θcos10θ?

OpenStudy (callisto):

Nope.. 2sinxcosx = sin2x

OpenStudy (anonymous):

how did cosx disappear?

OpenStudy (callisto):

It's an identity.... sin 2x = sin (x+x) = sinx cosx + cosx sinx = 2 sinx cosx

OpenStudy (anonymous):

i get the identity part. but is it possible to show me in steps after 2(sin5θcos5θ)?

OpenStudy (callisto):

As I've told you, put x=5θ into the identity sin2x = 2sinxcosx For the right, you get something familiar, which is a part of your question For the left, it's something you want. What do you get for the left?

OpenStudy (anonymous):

sin10θ ?

OpenStudy (callisto):

Yes. But for the question, you have 4 sin5θ cos5θ So, you'll have 2 (2sin5θ cos 5θ) ^ Can you simplify it?

OpenStudy (anonymous):

I dont know how to simplify this....

OpenStudy (callisto):

1. simplify this: (2sin5θ cos 5θ) 2. Multiply the result you get by 2 What do you get for the first step?

OpenStudy (anonymous):

1) sin2a

OpenStudy (anonymous):

i mean sin25θ

OpenStudy (callisto):

Why is it 25θ??

OpenStudy (anonymous):

so 4sin5θ cos5θ = sin2x and x=5θ so sin2(5θ)=sin10θ ...isn't it done?

OpenStudy (callisto):

50% correct. 4sinx cosx = 2sin2x Put x = 5θ into the above identity.

OpenStudy (anonymous):

Oh because 2sinx cos x = sin2x and here, its 4sinx cos x so we put 2sin2x??

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

Ah thank you very much!! makes much more sense

OpenStudy (anonymous):

Is there such a question that it can be like (any number)sinxcosx?

OpenStudy (callisto):

I'm not sure... but 2sinxcosx = sin2x If you have sinxcosx, you can rewrite it as 1/2 (sin2x) Btw, remember to put x = 5θ back to your question!

OpenStudy (anonymous):

so if it was sinxcosx= 1/2sinx?

OpenStudy (callisto):

No! sinxcosx= 1/2sin2x

OpenStudy (callisto):

It's double angle formula!

OpenStudy (anonymous):

oh Yeah i made a mistake, i meant by 1/2sin2x

OpenStudy (callisto):

Yes. \(sinxcosx = \frac{1}{2}sin2x\)

OpenStudy (anonymous):

thank you!!

OpenStudy (callisto):

Welcome!